# ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People

### v1

@@ 6. Discussion
In this section, we summarize our key findings to address our research questions, discuss potential use cases for the object-search system, and reflect on requirements to be addressed by future work for object-search systems to be of use to people who are blind.
RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
Overall, our experience with the design and implementation of ObjectFinder suggests that it is feasible to create assistive technology for object search that combines detection-based and description-based technologies, and that can be leveraged by blind people wishing to examine unfamiliar environments. In particular, our work shows that such a system can facilitate exploration of different settings (i.e., living and office environments), and that it can be a first step toward building a mental map of an unknown space. In particular, the description part of the ObjectFinder prototype offered an overview that helped participants understand the environment, know what to expect in the scenario, and discover unknown objects for subsequent searches. Compared to the description-only and detection-only applications, systems with the design of ObjectFinder have potential to provide a guiding function and the comprehensive information necessary for object search: localization information (distance and direction) and the relationship among the surrounding objects. However, our findings also show that there are several design challenges that remain unaddressed. For example, open questions regarding hardware design and interaction features remain, and addressing these is crucial when creating technology that people are interested in integrating into their daily lives. With respect to the information provided by the system, results suggest that the amount of information and whether to make additional details optional for object search, such as color and alert information, needs to be fine-tuned based on participants’ information processing strategies.
RQ2: What are users’ experiences and preferences regarding the open-vocabulary interactive object-search system, specifically comparing the description-based and the detection-based, with the hybrid ObjectFinder prototype?
Given the absence of existing technology to assist blind individuals in object search, most participants expressed enthusiasm for this open-vocabulary interactive object-search prototype. Participant ratings reflect satisfaction with the prototypical integration of functions within ObjectFinder, and suggest that cognitive load during system use was acceptable. In comparison with exclusively description-based or detection-based systems, our prototype effectively defines the target object, locates regions of interest, and delivers essential localization information and route planning, with interview results showing that this was appreciated by users. With respect to user experience with specific features of ObjectFinder, our work shows that detection of small objects was well-supported, whetheras the user experience when exploring larger space was less smooth, with open questions regarding user safety remaining (i.e., better obstacle recognition).
6.1. Future Directions for the Development of Object-Search Systems
Our research highlights that object-search systems have potential as an assistive technology for people who are blind. Here, we want to highlight future directions for the development of such systems, outlining challenges and opportunity for the Human-Computer Interaction and accessibility research communities.
6.1.1. Potential Features for Future Integration
@@
Providing better and more tailored descriptions. Providing descriptive overviews of unfamiliar environments is essential for revealing unexpected objects and should be relative to the user’s position. Based on our findings, we recommend to include both distance and direction in descriptions, with far objects quantified in meters and nearer objects described in steps or as ‘right in front of you’. These details can facilitate room exploration, help in building mental maps, and assist in orientation.
Understanding advantages and drawbacks of additional information. Although the essential information for object search, localization information and relationships among surrounding objects, is well-defined, participants had mixed feelings about additional details like color and alert information. Consequently, the future object-search system should allow users to select the amount of information by incorporating options to skip information or to stop and continue information output. Likewise, future work should explore user preferences for the types of information to be included, and implications of additional unrequested descriptions for user experience and aspects such as cognitive load.
@@
6.1.2. Tensions and Concerns Regarding Vision- and AI-based Assistive Technology
@@
There are tensions and concerns that need to be resolved for systems such as ObjectFinder effectively and safely supporting object finding.
@@
User habits and needs regarding lighting conditions for camera-based systems. With respect to technical requirements of current approaches to computer vision systems, we note that lighting conditions are crucial for camera-based systems. However, this conflicts with the fact that light plays a different role in the lives of blind people: Legally blind persons may not use light in their homes and workplaces in the same way as (typically) sighted system developers would anticipate  (Gonzalez Penuela et al., 2024), and people who do have residual vision may not find lighting conditions required by camera systems comfortable.
@@
Addressing safety concerns in the context of AI. Likewise, we noted instances in which ObjectFinder did not recognize furniture, e.g., when living room furnishings were lower than those in office environments, and not in view of the camera worn by participants. While there are established strategies for people who are blind or have low vision to use gaze for environmental scanning (Riddering, 2023), no specific scanning strategies for smart glasses tailored to blind individuals currently exist. With advancements in wearable systems and smart glasses, researching and integrating these specific scanning techniques into mobility training is both viable and beneficial. Moreover, incorporating sensors and cameras with wider angles, like LiDAR (Liu et al., 2021a) and omnidirectional camera (Kawaharazuka et al., 2024), could enhance scene perception over larger areas, reducing the need for physical scanning efforts. However, despite potential technological solutions (which may come with new challenges), this specific instance highlights tensions around safety: On the one hand, users are invited to rely on systems for object detection, on the other hand, it is known that vision- and AI-based systems can be unreliable in specific situations (e.g., in the context of autonomous driving (Cummings and Bauchwitz, 2024)), and even more so in the context of disability (e.g., see (Kane et al., 2020)). Thus, there remains a tension between what AI-based assistive technology seeks to offer, and what it can realistically provide, which is an aspect that needs to be communicated with nuance, and should be negotiated together with target audiences in the context of future work.
@@
Understanding the limitations of technology for object search. Finally, the Human-Computer Interaction and accessibility communities have previously discussed issues surrounding technologies focusing on independence of users. In particular, Vincenzi (Vincenzi, 2021) contributed a critical appraisal of assistive technology for navigation of blind people, suggesting that there were instances where working with other people was more relevant than applying a technology-based solution. In this context, the principle of interdependence (Bennett et al., 2018) is relevant, i.e., the fact that we all exist within relationship with our environment, and that assistive technology should not only consider the individual user but also the (social) context within which it exists, and implications of its design under consideration of opportunities to create collective access. Here, we need to ask critical questions around specific features of ObjectFinder, and we want to leave you with one example: Is it really necessary for a blind person to use the system in the workplace, or could non-disabled colleagues make a bigger effort to not misplace or alter their desk?
7. Limitations and Future Work
There are a few limitations that need to be taken into account when interpreting our findings. This research, conducted in small indoor settings of our lab, does not investigate the effectiveness of ObjectFinder in larger, real-world public, private, or outdoor environments. We therefore reached only participants from nearby regions. The requirement for a WiFi connection by ObjectFinder, due to its use of GPT-4, limits its operational range and may introduce security concerns akin to those observed with Be My Eyes. Although ObjectFinder possesses MLLM capabilities to infer surroundings beyond the immediate view, it lacks the ability to precisely remember explored spaces, rendering every space unfamiliar to its current version and thus providing known information to the user. Additionally, while participants preferred our ObjectFinder when paired with efficient hardware (Waisberg et al., 2024), we only had them try the hardware and imagine its implementation, as it was not actually deployed on more efficient hardware.
We see several avenues for future work. With the development of smart glasses, we will pursue collaborations with companies to create a more comfortable and affordable system, which will integrate ObjectFinder’s functionality to support object search and build up mental maps. After addressing the challenges and implementing the suggestions from participants, we plan to conduct a larger-scale quantitative evaluation of the system with blind individuals in real private and public scenarios across different cities. This approach would provide an opportunity to engage a larger number of participants and to evaluate the effectiveness of ObjectFinder in real-world scenarios, which we initially opted not to do at scale. Our first goal was to assess whether the open-vocabulary interactive object-search system addresses existing challenges in daily life and whether our implementation delivers an independent and confident experience to the target group in controlled settings.
@@ 8. Conclusion
In this work, we explored the design and development of a prototype that combines detection and description to enable open-vocabulary interactive object search for blind people. With our prototype, we address shortcomings of existing systems that are either description- or detection-based: locating regions of interest and discovering incidental targets. Results of the exploratory user study suggest that the approach is promising, offering blind people a tool to develop a better understanding of unfamiliar spaces. Overall, our work represents an initial step toward developing AI-based assistive technology that supports object search, providing first insights into user requirements and application challenges. Here, we hope that our work will encourage and facilitate further development of object-search systems, and that it will inspire future studies into the experiences that blind people have with such technologies.

### v2

@@ 6. Discussion
Our research highlights that object-search systems have potential as an assistive technology for people who are blind. Here, we want to discuss future directions for the development of such systems, outlining challenges and opportunity for the Human-Computer Interaction and accessibility research communities.
6.1. Potential Features for Future Integration
@@
Providing better and more tailored descriptions. Providing descriptive overviews of unfamiliar environments is essential for revealing unexpected objects and should be relative to the user’s position. Based on our findings, we recommend to include both distance and direction in descriptions, with far objects quantified in meters and nearer objects described in steps or as “right in front of you”. These details can facilitate room exploration, help in building mental maps, and assist in orientation.
Understanding advantages and drawbacks of additional information. Although the essential information for object search, localization information, and relationships among surrounding objects, is well-defined, participants had mixed feelings about additional details like color and alert information. Consequently, the future object-search system should allow users to select the amount of information by incorporating options to skip information or to stop and continue information output. Likewise, future work should explore user preferences for the types of information to be included, and implications of additional unrequested descriptions for user experience and aspects such as cognitive load.
@@
6.2. Tensions and Concerns Regarding Vision- and AI-based Assistive Technology
@@
There are tensions and concerns that need to be resolved for systems such as ObjectFinder to effectively and safely support object searching.
@@
User habits and needs regarding lighting conditions for camera-based systems. With respect to the technical requirements of current approaches to computer vision systems, we note that lighting conditions are crucial for camera-based systems. However, this conflicts with the fact that light plays a different role in the lives of blind people: Legally blind persons may not use light in their homes and workplaces in the same way as (typically) sighted system developers would anticipate  (Gonzalez Penuela et al., 2024), and people who do have residual vision may not find lighting conditions required by camera systems comfortable.
@@
Addressing safety concerns in the context of AI. Likewise, we noted instances in which ObjectFinder did not recognize furniture, e.g., when living room furnishings were lower than those in office environments, and not in view of the camera worn by participants. While there are established strategies for people who are blind or have low vision to use gaze for environmental scanning (Riddering, 2023), no specific scanning strategies for smart glasses tailored to blind individuals currently exist. With advancements in wearable systems and smart glasses, researching and integrating these specific scanning techniques into mobility training is both viable and beneficial. Moreover, incorporating sensors and cameras with wider angles, like LiDAR (Liu et al., 2021) and omnidirectional camera (Kawaharazuka et al., 2024), could enhance scene perception over larger areas, reducing the need for physical scanning efforts. However, despite potential technological solutions (which may come with new challenges), this specific instance highlights tensions around safety: On the one hand, users are invited to rely on systems for object detection, on the other hand, it is known that vision- and AI-based systems can be unreliable in specific situations (e.g., in the context of autonomous driving (Cummings and Bauchwitz, 2024)), and even more so in the context of disability (e.g., see (Kane et al., 2020)). Thus, there remains a tension between what AI-based assistive technology seeks to offer, and what it can realistically provide, which is an aspect that needs to be communicated with nuance, and should be negotiated together with target audiences in the context of future work.
@@
Understanding the limitations of technology for object search. Finally, the Human-Computer Interaction and accessibility communities have previously discussed issues surrounding technologies focusing on the independence of users. In particular, Vincenzi (Vincenzi, 2021) contributed a critical appraisal of assistive technology for navigation of blind people, suggesting that there were instances where working with other people was more relevant than applying a technology-based solution. In this context, the principle of interdependence (Bennett et al., 2018) is relevant, i.e., the fact that we all exist within relationship with our environment, and that assistive technology should not only consider the individual user but also the (social) context within which it exists, and implications of its design under consideration of opportunities to create collective access. Here, we need to ask critical questions around specific features of ObjectFinder, and we want to leave you with one example: Is it really necessary for a blind person to use the system in the workplace, or could non-disabled colleagues make a bigger effort to not misplace or alter their desk?
@@ 7. Conclusion
In this work, we explored the design and development of a prototype that combines detection and description to enable open-vocabulary interactive object search for blind people. With our prototype, we address shortcomings of existing systems that are either description- or detection-based: locating regions of interest and discovering incidental targets. The system feedback is tailored to various user intents, and our exploratory user study suggests that this approach is promising, as it provides essential egocentric localization and allocentric scene context while enabling interactive object search. Overall, our work represents an initial step towards developing AI-based assistive technology that supports object search, providing first insights into user requirements and application challenges. Here, we hope that our work will encourage and facilitate further development of object-search systems, and that it will inspire future studies into the experiences that blind people have with such technologies.